OpenStudy (anonymous):

A model rocket is launched with an initial upward velocity of 175 ft/s The rocket's height h (in feet) after t seconds is given by the following. h = 175t - 16t^2 Find all values of t for which the rocket's height is 85 feet. Round your answer(s) to the nearest hundredth.

OpenStudy (aihberkhan):

Okay. It is quite simple actually. The trick is to pull out all the numbers and realize why they are there. So plug in all the values: \(85 = 175t - 16t^2\) Now simply continue to solve: \(16t^2 - 175t + 85 = 0\) \(t = (175±√((-175)2 - 4(16)(85))) / 2(16)\) \(t = (175±158.70) / 32\) \(t = 10.43\) seconds \(t = 0.51 \) seconds

OpenStudy (aihberkhan):

That's it! So the final answer is \(t = 0.51 \) seconds @theyenvydiamond

OpenStudy (aihberkhan):

Hope this helped! Have a great day! :) A medal would also be appreciated! Just click best response next to my name! Also, a fan would be appreciated as well! Just hover over my icon and click "Become A Fan". This will allow you to see every time I am online! :) If you see that I am online and need help with a question, just tag me in your question! @theyenvydiamond

OpenStudy (anonymous):

Thank you so much! :)